### v1

# 3rd Generation Partnership Project 5G Fifth Generation

10G-EPON
10 Gigabit EPON
3GPP
3rd Generation Partnership Project
5G
Fifth Generation
6G
Sixth Generation
A-CPI
A Controller Plane Interface
A-RoF
Analog Radio-over-Fiber
ABNO
Application-Based Network Operations
ADC
Analog-to-Digital Converter
AE
Allocative Efficiency
amc
adaptive modulation and coding
AON
Active Optical Network
AI
Artificial Intelligence
ML
Machine Learning
API
Application Programming Interface
AWG
Arrayed Waveguide Grating
B2B
Business to Business
B2C
Business to Consumer
B-RAS
Broadband Remote Access Servers
BaaS
Blockchain as a Service
BB
Budget Balance
BBF
BroadBand Forum
BBU
Base Band Unit
BER
Bit Error Rate
BEREC
Body of European Regulators for Electronic Communications
BF
beamforming
BFT
Byzentine fault tolerant
BGP
Border Gateway Protocol
BMap
Bandwidth Map
BPM
Business Process Management
BS
Base Station
BSC
Base Station Controller
BtB
Back to Back
BW
Bandwidth
C2C
Consumer to Consumer
C-RAN
Cloud Radio Access Network
C-RoFN
Cloud-based Radio over optical Fiber Network
CA
Certificate Authority
CapEx
Capital Expenditure
CDMA
code division multiple access
CDR
Clock Data Recovery
CFO
carrier frequency offset
CIR
Committed Information Rate
CO
Central Office
CoMP
Coordinated Multipoint
COP
Control Orchestration Protocol
CORD
Central Office Rearchitected as a Data Centre
CPE
Customer Premises Equipment
CPRI
Common Public Radio Interface
CPU
Central Processing Unit
CMC
Conflict Mitigation Controller
CMS
Conflict Management System
CQI
channel quality indicator
CR
cognitive radio
CS
Central Station
CSI
channel state information
D2D
Device to Device
D-CPI
D Controller Plane Interface
D-RoF
Digital Radio-over-Fiber
DAC
Digital-to-Analog Converter
DAS
distributed antenna system
DAS
Distributed Antenna Systems
DBA
Dynamic Bandwidth Allocation
DBRu
Dynamic Bandwidth Report upstream
DC
Direct Current
DL
Downlink
DLT
Distributed Ledger Technology
DMT
Discrete Multitone
DNS
Domain Name Server
DPDK
Data Plane Development Kit
DSB
Double Side Band
DSL
Digital Subscriber Line
DSLAM
Digital Subscriber Line Access Multiplexer
DSP
Digital Signal Processing
DSS
Distributed Synchronization Service
DU
Distributed Unit
DUE
D2D user equipment
DWDM
Dense Wave Division Multiplexing
E-CORD
Enterprise CORD
EAL
Environment Abstraction Layer
EDF
Erbium-Doped Fiber
EFM
Ethernet in the First Mile
EMBS
Elastic Mobile Broadband Service
EON
Elastic Optical Network
EPC
Evolved Packet Core
EPON
Ethernet Passive Optical Network
eMBB
enhanced Mobile Broadband
ETSI
European Telecommunications Standards Institute
EVM
Error Vector Magnitude
FANS
Fixed Access Network Sharing
FASA
Flexible Access System Architecture
FBMC
filter bank multi-carrier
FBMC
Filterbank Multicarrier
FCC
Federal Communications Commission
FDD
Frequency Division Duplex
FDMA
frequency-division multiple access
FEC
Forward Error Correction
FFR
Fractional Frequency Reuse
FFT
Fast Fourier Transform
FPGA
Field-Programmable Gate Array
FSO
Free-Space-Optics
FSR
Free Spectral Range
FTN
faster-than-nyquist
FTTC
Fiber-to-the-Curb
FTTH
Fiber-to-the-Home
FTTx
Fiber-to-the-x
G.Fast
Fast Access to Subscriber Terminals
GEM
G-PON encapsulation method
GFDM
Generalized Frequency Division Multiplexing
GMPLS
Generalized Multi-Protocol Label Switching
GPON
Gigabit Passive Optical Network
GPP
General Purpose Processor
GPRS
General Packet Radio Service
GSM
Global System for Mobile Communications
GTP
GPRS Tunneling Protocol
GUI
Graphical User Interface
HA
Hardware Accelerator
HARQ
Hybrid-Automatic Repeat Request
HD
half duplex
HetNet
heterogeneous networks
IaaS
Infrastructure as a Service
I-CPI
I Controller Plane Interface
I2RS
Interface 2 Routing System
IA
Interference Alignment
IAM
Identity and Access Management
IBFD
in-band full duplex
IC
Incentive Compatibility
ICIC
inter-cell interference coordination
ICT
information and communications technology
IEEE
Institute of Electrical and Electronics Engineers
IETF
Internet Engineering Task Force
IF
Intermediate Frequency
IFFT
inverse FFT
ITS
Intelligent Transport Systems
InP
Infrastructure Provider
IoT
Internet of Things
IP
Internet Protocol
IQ
In-phase/Quadrature
IR
Individual Rationality
IRC
Interference Rejection Combining
ISI
inter-symbol interference
ISP
Internet Service Provider
BFT
Byzantine-Fault-Tolerant
IV
Intelligent Vehicle
ITU
International Telecommunication Union
KPI
Key Performance Indicator
KVM
Kernel-based Virtual Machine
L2
Layer-2
L3
Layer-3
LAN
Local Area Network
LO
Local Oscillator
LOS
Line Of Sight
LR-PON
Long Reach PON
LSP-DB
Label Switched Path Database
LTE-A
Long Term Evolution Advanced
LTE
Long Term Evolution
M-CORD
Mobile CORD
M-MIMO
massive MIMO
M2M
machine-to-machine
MAC
Medium Access Control
ME
Merging Engine
MIMO
multiple input multiple output
MME
Mobility Management Entity
mmWave
millimeter wave
MNO
mobile network operator
MPLS
Multiprotocol Label Switching
MRC
Maximum Ratio Combining
MSC
Mobile Switching Centre
MSP
Membership Service Providers
MSR
Multi-Stratum Resources
MTC
machine type communication
mMTC
massive Machine Type Communications
MVNO
Mobile Virtual Network Operator
MVNP
mobile virtual network provider
MZI
Mach-Zehnder Interferometer
MZM
Mach-Zehnder Modulator
NE
Network Element
NFV
Network Function Virtualization
NFVaaS
Network Function Virtualization as a Service
NG-PON2
Next-Generation Passive Optical Network 2
NGA
Next Generation Access
NLOS
None Line Of Sight
NMS
Network Management System
NRZ
Non Return-to-Zero
NTT
Nippon Telegraph and Telephone
OBPF
Optical BandPass Filter
OBSAI
Open Base Station Architecture Initiative
ODN
Optical Distribution Network
OFC
Optical Frequency Comb
OFDM
orthogonal frequency-division multiplexing
OFDMA
orthogonal frequency-division multiple access
OLO
Other Licensed Operator
OLT
Optical Line Terminal
ONF
Open Networking Foundation
ONU
Optical Network Unit
OOB
out-of-band
OOK
On-off Keying
OpenCord
Central Office Re-Architected as a Data Center
OpEx
Operating Expenditure
OSI
Open Systems Interconnection
OSS
Operations Support Systems
OTT
Over-the-Top
OXM
OpenFlow Extensible Match
P2MP
Ethernet over point-to-multipoint
P2P
Point-to-Point
PAM
Pulse Amplitude Modulation
PAPR
peak-to-average power rating
PAPR
Peak-to-Average Power Ratio
PBMA
Priority Based Merging Algorithm
PCE
Path Computation Elements
PCEP
Path Computation Element Protocol
PCF
Photonic Crystal Fiber
PD
Photodiode
PDCP
RD Control Protocol
PDF
probability distribution function
PGW
Packet Gateway
PHY
physical Layer
PIR
Peak Information Rate
PMD
Polarization Division Multiplexing
PON
Passive Optical Network
POTS
Plain Old Telephone Service
PoET
Proof of Elapsed Time
PoW
Proof of Work
PoS
Proof of Stake
PTP
Precision Time Protocol
PWM
Pulse Width Modulation
QAM
Quadrature Amplitude Modulation
QoE
Quality of Experience
QoS
Quality of Service
QPSK
Quadrature Phase Shift Keying
R-CORD
Residential CORD
RA
Resource Allocation
RAN
Radio Access Network
RAT
Radio Access Technology
RB
resource block
RFIC
radio frequency integrated circuit
RIC
RAN Intelligent Controller
RLC
Radio Link Control
RN
Remote Node
ROADM
Reconfigurable Optical Add Drop Multiplexer
RoF
Radio-over-Fiber
RRC
Radio Resource Control
RRH
Remote Radio Head
RRPH
Remote Radio and PHY Head
RRU
Remote Radio Unit
RSOA
Reflective Semiconductor Optical Amplifier
RU
Remote Unit
Rx
receiver
SC-FDMA
single carrier frequency division multiple access
SCM
single carrier modulation
SD-RAN
Software Defined Radio Access Network
SDAN
Software Defined Access Network
SDMA
Semi-Distributed Mobility Anchoring
SDN
Software Defined Networking
SDR
Software Defined Radio
SFBD
Single Fiber Bi-Direction
SGW
Serving Gateway
SI
self-interference
SIC
self-interference cancellation
SIMO
Single Input Multiple Output
SLA
Service Level Agreement
SMF
Single Mode Fiber
SNMP
Simple Network Management Protocol
SNR
Signal-to-Noise Ratio
SP
Service Providers
Split-PHY
Split Physical Layer
SRI-OV
Single Root Input/Output Virtualization
SU
secondary user
SUT
System Under Test
T-CONT
Transmission Container
TC
Transmission Convergence
TCO
Total Cost of Ownership
TD-LTE
Time Division LTE
TDD
Time Division Duplex
TDM
Time Division Multiplexing
TDMA
Time Division Multiple Access
TED
Traffic Engineering Database
TEID
Tunnel endpoint identifier
TLS
Transport Layer Security
TPS
Transactions per Second
TTI
transmission time interval
TWDM
Time and Wavelength Division Multiplexing
Tx
transmitter
UD-CRAN
Ultra-Dense Cloud Radio Access Network
UDP
User Datagram Protocol
UE
User Equipment
UF-OFDM
Universally Filtered OFDM
UFMC
Universally Filtered Multicarrier
UL
Uplink
UMTS
Universal Mobile Telecommunications System
URLLC
Ultra-Reliable Low-Latency Communication
USRP
Universal Software Radio Peripheral
V2I
Vehicle to Infrastructure
V2V
Vehicle to Vehicle
vBBU
virtualized BBU
vBMap
Virtual Bandwidth Map
vBS
virtual Base station
VCG
Vickery-Clarke-Groves
vCPE
virtual CPE
vCPU
Virtual Central Processing Unit
vDBA
virtual DBA
VDSL2
Very-high-bit-rate digital subscriber line 2
VLAN
Virtual Local Area Network
VM
Virtual Machine
VNF
Virtual Network Function
VNI
visual networking index
VNO
Virtual Network Operator
VNTM
Virtual Network Topology Manager
vOLT
virtual OLT
VPE
virtual Provider Edge
VPN
Virtual Private Network
V2X
Vehicle to Everything
VTN
Virtual Tenant Network
VULA
Virtual Unblunded Local Access
WAN
Wide Area Network
WAP
Wireless Access Point
WCDMA
wide-band code division multiple access
WDM-PON
Wavelength Division Multiplexing - Passive Optical Network
WDM
Wavelength Division Multiplexing
WiMAX
Worldwide Interoperability for Microwave Access
WRPR
Wired-to-RF Power Ratio
XG-PON
10 Gigabit PON
XGEM
XG-PON encapsulation method
XOS
XaaS Operating System
RIC
RAN Intelligent Controller
Near-RT-RIC
Near Real Time RAN Intelligent Controller
Non-RT-RIC
Non Real Time RAN Intelligent Controller
xApp
Extended Application
rApp
Remote Application
CMS
Conflict Mitigation System
EG
Eisenberg-Galle
NSWF
Nash’s Social Welfare Function
CMC
Conflict Mitigation Controller
CDC
Conflict Detection Controller
PMon
Performance Monitoring
RCP
Recently Changed Parameter
PGD
Parameter Group Definition
RCPG
Recently Changed Parameter Group
DCKD
Decision Correlated with KPI Degradation
KDO
KPI Degradation Occurrences
MNO
Mobile Network Operator
SDL
Shared Data Layer
SON
Self Organising Network
SF
SON Function
AM
Arithmetic Mean
PKR
Parameter and KPI Ranges
KPI
Key Performance Indicator
ICP
Input Control Parameter
RF
Radio Frequency
C-RAN
Cloud Radio Access Network
MARL
Multi Agent Reinforcement Learning
CCO
Capacity and Coverage Optimization
ES
Energy Saving
MRO
Mobility Robustness Optimization
MLB
Mobility Load Balancing
V-RAN
Virtualized Radio Access Network
O-RAN
Open Radio Access Network
D-RAN
Distributed Radio Access Network
5GB
5G and Beyond
O-RU
Open Radio Unit
O-DU
Open Distribution Unit
O-CU
Open Centralised Unit
O-CU-CP
Open Centralised Unit - Control Plane
O-CU-UP
Open Centralised Unit - User Plane
D-SON
Distributed SON
C-SON
Centralised SON
CIO
Cell Individual Offset
CBR
Call Block Ratio
HOR
Handover Ratio
CAN
Cognitive Autonomous Network
CF
Cognitive Function
QACM
QoS-Aware Conflict Mitigation
ANN
Artificial Neural Network
PR
Polynomial Regression
MSE
Mean Squared Error
EVS
Explained Variance Score
CS
Conflict Supervision
TXP
Tx Power
TTT
Time-To-Trigger
CIO
Cell Individual Offset
NL
Neighbour List
RET
Radio Electrical Tilt
HYS
Handover Hysteresis
QACMP
QACM Priority
DRL
Deep Reinforcement Learning
gNB
Next-generation NodeB
RSRP
Reference Signal Received Power
This version of the paper has been submitted for publication to IEEE and may be removed upon request or copyright issues. xApp-Level Conflict Mitigation in O-RAN, a Mobility Driven Energy Saving Case
Abdul Wadud12,  Fatemeh Golpayegani1,  and Nima Afraz 1 Corresponding author: Abdul Wadud (email: abdul.wadud@ucdconnect.ie). 1School of Computer Science, University College Dublin, Ireland 2Bangladesh Institute of Governance and Management, Dhaka, Bangladesh
Abstract
This paper investigates the emerging challenges of conflict detection and mitigation in Open Radio Access Network (O-RAN). Conflicts between xApps can arise that affect network performance and stability due to the disaggregated nature of O-RAN. This work provides a detailed theoretical framework of Extended Application (xApp)-level conflicts, i.e., direct, indirect, and implicit conflicts. Leveraging conflict graphs, we further highlight how conflicts impact Key Performance Indicators and explore strategies for conflict detection using Service Level Agreements and Quality of Service (QoS) thresholds. We evaluate the effectiveness of several mitigation strategies in a simulated environment with Mobility Robustness Optimization (MRO) and Energy Saving (ES) xApps and present experimental results showing comparisons among these strategies. The findings of this research provide significant insights for enhancing O-RAN deployments with flexible and efficient conflict management.
Index Terms:
O-RAN, Conflict, SLA, QoS, MRO, ES, xApp, and KPI.
I Introduction
The introduction of O-RAN architecture has promised enhanced flexibility and interoperability in the telecommunications industry. However, this disaggregated approach of integrating components from multiple vendors presents a significant challenge of conflict management. While the traditional Radio Access Network (RAN) relied on single-vendor solutions that come with in-house vendor spesific conflict resolution strategies, O-RAN’s multi-vendor nature needs a robust conflict mitigation strategy to cope with conflicts between different vendors’ components. Conflicts arise when components compete for shared resources or set conflicting configurations that control the same network environment. This can lead to performance degradation, instability, and security vulnerabilities in the network. Therefore, handling these conflicts are now a priority and a major obstcle for wide deployment on O-RAN. The following provides an initial understanding of different types of conflicts in O-RAN based on the state-of-the-art.
I-A Categorization of Conflicts:
Conflicts are categorized into horizontal and vertical conflicts, which are further classified into intra-component and inter-component conflicts [1, 2, 3].
Horizontal conflicts occur between components at the same hierarchical level, for instance, between two xApps (applications for near-real-time control) within a Near Real Time RAN Intelligent Controller (Near-RT-RIC) [2]. It is further classified into intra-component and inter-component conflicts:
  • Intra-component conflicts occur within a component, like conflicting configurations between xApps in a Near-RT-RIC [2, 3]. These are further classified into three types:
    • Direct Conflict: These occur when two or more xApps attempt to simultaneously control the same network parameter [1, 4, 2, 5, 6]. For instance, one xApp might try to increase the transmission power for a specific cell, while another aims to decrease it for energy saving. Direct conflicts are generally easier to detect since the conflicting parameter is directly observable [7, 3]. However, simply prioritizing one xApp’s preference over another might not be the optimal solution. A more effective approach would involve finding a configuration that balances these conflicting objectives, potentially maximizing overall network utility [2].
    • Indirect Conflicts: These arise when different xApps control distinct parameters that ultimately influence the KPI [1, 6, 8, 2, 9]. For example, one xApp might modify the cell individual offset (CIO) to balance the load across cells, while another adjusts antenna tilts. Both actions indirectly impact the cell boundaries and handover decisions, potentially causing conflicts [10, 8]. Detecting indirect conflicts necessitates post-action analysis, observing how different xApp actions collectively impact the relevant KPI. This is more challenging than detecting direct conflicts as the relationship between parameter changes and KPI impacts might not be immediately apparent [3].
    • Implicit Conflicts: These present the most significant challenge for detection and mitigation [1, 7]. Implicit conflicts occur when the actions of multiple xApps, while individually aligned with their specific objectives, result in an undesirable overall network state [4, 8]. This often involves subtle interactions between xApps with different goals, making the source of the conflict difficult to pinpoint. For example, an xApp trying to maximize quality of service (QoS) for a group of users might interfere with another xApp aiming to minimize handover rates, leading to unexpected performance degradation. Identifying and resolving implicit conflicts often requires sophisticated monitoring and analysis of network behavior over time [1].
  • Inter-component conflicts involve components from different areas, such as conflicting decisions from xApps in neighboring Near-RT-RICs [3, 2].
Vertical conflicts involve components from different hierarchical levels, such as a conflict between a Near-RT-RIC and a Non-RT RIC [2, 3].
Addressing various conflicts requires a multi-pronged approach. Existing literature suggests several strategies:
I-A1 Conflict Detection and Mitigation Frameworks:
Adamczyk and Kliks [1] propose a conflict mitigation framework (CMF) within the Near-RT-RIC. The framework focuses on detecting direct, indirect, and implicit conflicts between xApps. Direct conflicts arise from consecutive, contradictory decisions affecting the same parameters. Indirect conflicts occur when xApps impact shared KPIs through different parameters. Implicit conflicts involve xApp actions aligning with individual objectives but contradicting overall network goals. A more detailed description of these conflicts are discussed in Section. II. The framework in [1] utilizes a Conflict Mitigation (CM) component in the Near-RT-RIC to identify and resolve these conflicts. However, current implementations primarily focus on specific xApps, like Mobility Robustness Optimization (MRO) and Mobility Load Balancing (MLB).
I-A2 Team Learning and Cooperative Approaches:
Zhang et al. in [11] suggest a team learning algorithm based on Deep Q-learning to encourage cooperation between xApps. By sharing information about their intended actions, xApps can make more informed decisions that benefit the overall network performance. This approach demonstrates promising results, leading to improved throughput and reduced packet drop rates. However, scalability remains a concern, as the complexity increases with the number of cooperating xApps.
I-A3 QoS-Aware Conflict Mitigation:
Recognizing the importance of maintaining Quality of Service (QoS), our previous work [3] introduces the QACM (QoS-Aware Conflict Mitigation) method. QACM considers the QoS benchmarks of individual xApps during conflict mitigation, ensuring their requirements are met. It utilizes game theory principles, specifically Nash’s Social Welfare Function (NSWF) and Eisenberg-Galle (EG) solutions, to find an optimal balance between conflicting parameters. Initial results indicate QACM’s effectiveness in upholding QoS thresholds compared to benchmark methods. However, further research is needed to develop more sophisticated KPI prediction models and test its practicality in real-world RAN deployments.
I-A4 Motivation and Contribution of this Research:
While significant progress has been made in understanding and mitigating conflicts in Open RAN, several challenges remain:
  • As Open RAN is a new concept, and there is limited knowledge about conflicts within it. The subtle difference between indirect and implicit conflicts makes reliably detecting and evaluating these potential issues challenging. Developing a low-latency detection mechanism is crucial to ensure the reliability and stability of O-RAN deployments.
  • Dynamic Detection and Mitigation: Existing solutions often focus on static conflict resolution. However, the dynamic nature of network traffic and user demands necessitates adaptive and real-time conflict detection and mitigation strategies.
  • Scalability and Complexity: As O-RAN networks expand in size and complexity, managing conflicts between numerous xApps and Remote Applications from diverse vendors will become increasingly challenging. Scalable and efficient conflict detection and mitigation mechanisms are essential for the large-scale adoption of Open RAN.
In light of these challenges, this research focuses on:
  • Providing a detailed theoretical framework to explain each type of xApp conflict that helps simplifying conflict detection and mitigation.
  • Utilizing conflict graphs for more clear representation and better understanding of potential conflict.
  • Identifying differences between conflict types systematically and developing a low-latency conflict detection framework adaptable to evolving network conditions.
  • Reviewing and analyzing the performance of state-of-the-art mitigation strategies.
The structure of the paper is as follows: Section II provides a theoretical modeling and evaluation of xApp conflicts, Section III presents a discussion on conflict detection and mitigation strategies, Section IV discusses simulation experiments and results with ES and MRO xApps, and the paper concludes in Section V.
II Understanding xApp Conflicts in Open RAN
II-A Theoretical Modeling of Intra-Component Conflicts
We develop a theoretical framework for modeling intra-component conflicts, focusing on the interactions between different xApps within the Near-RT-RIC. This framework uses conflict graphs to represent potential conflict points and their characteristics. Let us consider, there are $n$ number of xApps present in a Near-RT-RIC, represented by $x_{n}\in X$. Additionally, there are $i$ numbers of Input Control Parameters and $j$ number of KPIs, each represented as $p_{i}\in P$ and $k_{j}\in K$, respectively. Each xApp has a set of ICPs denoted by $I_{x_{n}}$ and a set of KPIs represented by $K_{x_{n}}$. Intra-component conflicts are of three types as stated above in Section I-A.
We consider an example model with five stochastic xApps in Fig. 1 to explain each of the three types of conflicts. We consider only five xApps, because, our main goal is to cover all three types of conflicts and represent them in the most reader-friendly way in a limited space.
Figure 1: An example model of conflict with five stochastic xApps [3].
In Fig. 1, the set of input control parameters and KPIs for each xApp is defined as follows: $I_{x_{1}}=\{p_{1},p_{2}\}$, $I_{x_{2}}=\{p_{1},p_{2},p_{3}\}$, $I_{x_{3}}=\{p_{1},p_{4}\}$, $I_{x_{4}}=\{p_{5},p_{6}\}$, $I_{x_{5}}=\{p_{7},p_{8}\}$, $K_{x_{1}}=\{k_{1}\}$, $K_{x_{2}}=\{k_{2}\}$, $K_{x_{3}}=\{k_{3}\}$, $K_{x_{4}}=\{k_{41},k_{42}\}$, and $K_{x_{5}}=\{k_{5}\}$. The set of all xApps considered in this study is: $X=\{x_{1},x_{2},x_{3},x_{4},x_{5}\}$. The set of ICPs and KPIs considered in this study is: $P=\{p_{1},p_{2},p_{3},p_{5},p_{5},p_{6},p_{7}\}$ and $K=\{k_{1},k_{2},k_{3},k_{5},k_{41},k_{42}\}$. The parameter group for each KPI is formulated based on the control parameters that influence the KPI. The Algorithm 1 and K-P graph in Fig. 3 helps in identifying these groups. For example: $P_{k_{1}}^{G}=\{p_{1},p_{2}\}$, $P_{k_{2}}^{G}=\{p_{1},p_{2},p_{3}\}$, $P_{k_{3}}^{G}=\{p_{1},p_{4}\}$, $P_{k_{5}}^{G}=\{p_{7},p_{8}\}$, $P_{k_{41}}^{G}=\{p_{2},p_{5},p_{6}\}$, $P_{k_{42}}^{G}=\{p_{2},p_{5},p_{6}\}$. To understand how parameter group for each KPI is formed, let us consider $k_{1}$. From the K-P graph in Fig. 3, we see that $p_{1}$ and $p_{2}$ are parameters that influence $k_{1}$. Therefore, the parameter group for $k_{1}$ is: $P_{k_{1}}^{G}=\{p_{1},p_{2}\}$.
II-A1 Direct Conflict
A direct conflict occurs when an intersection operation between the set of input control parameters of two or more xApps results in a non-null set. This indicates that these xApps have conflicting input parameters. Based on the X-P graph in Fig. 2, specific direct conflict cases can be identified as follows:
  • $I_{x_{1}}\cap I_{x_{2}}=\{p_{1},p_{2}\}$: There is a direct conflict between $x_{1}$ and $x_{2}$ over the parameters $p_{1}$ and $p_{2}$.
  • $I_{x_{1}}\cap I_{x_{3}}=\{p_{1}\}$: There is a direct conflict between $x_{1}$ and $x_{3}$ over the parameter $p_{1}$.
  • $I_{x_{2}}\cap I_{x_{3}}=\{p_{1}\}$: There is a direct conflict between $x_{2}$ and $x_{3}$ over the parameter $p_{1}$.
II-A2 Indirect Conflict
An indirect conflict occurs when a group of input control parameters, belonging to different xApps, are associated with a common KPI. This type of conflict can be identified using the K-P graph. For instance:
  • $P_{k_{41}}^{G}=\{p_{2},p_{5},p_{6}\}$ and $P_{k_{42}}^{G}=\{p_{2},p_{5},p_{6}\}$: Here, $p_{2}\notin I_{x_{4}}$, but it still affects $k_{41}$ and $k_{42}$. Since $p_{2}\in I_{x_{1}}$ and $I_{x_{2}}$, we can say there is an indirect conflict of these xApps with $x_{4}$ over $p_{2}$.
II-A3 Implicit Conflict
An implicit conflict is identified when a KPI is affected by a parameter not originally considered part of its control parameter group. For example:
  • $P_{k_{5}}^{G}=\{p_{7},p_{8}\}$: This parameter group indicates that only modifying $p_{7}$ and $p_{8}$ should affect $k_{5}$. However, if during the RAN operation, $k_{5}$ gets affected by changing $p_{1}$, then we can say there is an implicit conflict between $x_{1},x_{2},x_{3}$ and $x_{5}$ over $p_{1}$ since $p_{1}\in I_{x_{1}}$, $I_{x_{2}}$, and $I_{x_{3}}$. Including $p_{1}$ inside $P_{k_{5}}^{G}$ changes it to $P_{k_{5}}^{G}=\{p_{1},p_{7},p_{8}\}$, making it an indirect conflict. Thus, an implicit conflict is a vague form of indirect conflict that can only be detected during RAN operation but can be modeled as an indirect conflict once detected.
Figure 2: X-P Graph with ICPs for xApps in Open RAN.
Figure 3: K-P Graph with KPIs for Parameters in Open RAN.
TABLE I: Comparison of Different xApps Conflict Types in Open RAN
Conflict Type Definition Mathematical Notation Example Conflict
Direct Conflict Occurs when two or more xApps share common input parameters $I_{x_{i}}\cap I_{x_{j}}\neq\emptyset$ $I_{x_{1}}\cap I_{x_{2}}=\{p_{1},p_{2}\}$
Indirect Conflict Occurs when different xApps’ input parameters influence the same KPI $P_{k_{i}}^{G}=\{p_{x},p_{y}\},P_{k_{j}}^{G}=\{p_{x},p_{y}\}$ $P_{k_{41}}^{G}=\{p_{2},p_{5},p_{6}\}$
Implicit Conflict Occurs when a KPI is influenced by parameters outside its control group $P_{k_{i}}^{G}=\{p_{x},p_{y}\}\Rightarrow p_{z}\text{ influences }k_{i}$ $P_{k_{5}}^{G}=\{p_{7},p_{8}\}\text{ but }p_{1}\text{ affects }k_{5}$
III Conflict Detection and Mitigation
Detecting conflicts is one of the most challenging aspects of conflict management in the Near-RT-RIC. It requires a deep understanding of each type of conflict and their distinctions. The comprehensive theoretical analysis presented in Section II and the comparison in Table I provide the foundation for a rule-based, low-latency conflict detection mechanism. xApp-level conflicts are typically parametric and can often be detected by tracing ICPs and monitoring KPI degradation events. Some KPIs are critical to maintain as per SLA agreements, while others are less so.
In the conflict management system proposed in [2, 3], the Mobile Network Operator (MNO) is responsible for defining the important KPIs that must be maintained at predefined thresholds. This paper adopts the Conflict Mitigation System (CMS) presented in our previous work [2, 3], which includes separate Conflict Detection Controller (CDC) and Conflict Mitigation Controller (CMC) components within the Near-RT-RIC. The CDC communicates back and forth with the Near-RT-RIC database while performing detection tasks.
The Near-RT-RIC database consists of several components, including Recently Changed Parameter (RCP), Parameter Group Definition (PGD), Recently Changed Parameter Group (RCPG), Parameter and KPI Ranges (PKR), Decision Correlated with KPI Degradation (DCKD), and KPI Degradation Occurrences (KDO). When a xApp is deployed in the Near-RT-RIC, the MNO is expected to provide the xApp details, including their ICPs ($I_{x_{n}}$) and KPIs ($K_{x_{n}}$). The PGD component of the database is equipped with Algorithm 1, which estimates the $P_{k}^{G}$ for all $k_{j}\in K_{x_{n}}$. The KDO component is responsible for storing and tracking KPI degradation events. If any SLA-sensitive KPI falls below the predefined threshold, it alerts the CDC of a possible conflict, along with the timestamp ($t_{clock}$) indicating when the event occured.
The RCP component tracks any parameter changes, aiding the CDC in identifying the exact parameter responsible for the conflict at $t_{clock}$. The step-by-step conflict detection process is illustrated in Fig. 4, which explains how a rule-based detection mechanism can identify specific types of conflicts.
Figure 4: Rule-based conflict detection mechanism.
Algorithm 1 Parameter Grouping Based on KPIs for xApps
0:  A set of xApps with their ICPs $I_{x_{n}}$ and KPIs $K_{x_{n}}$.
0:  Parameter group $P_{k}^{G}$ for $k_{j}$.
1:  Initialize a dictionary $param\_to\_kpis$ to map each ICP to its associated KPIs.
2:  Initialize a dictionary $P_{k}^{G}$ to map each KPI to its associated ICPs.
3:  for each xApp $x_{n}$ do
4:     for each ICP $p_{i}\in I_{x_{n}}$ do
5:        for each KPI $k_{j}\in K_{x_{n}}$ do
6:           Add $k_{j}$ to the list of KPIs for $p_{i}$ in $param\_to\_kpis$.
7:  for each pair of ICPs $(p_{i},p_{j})$ do
8:     Find the common KPIs between $p_{i}$ and $p_{j}$ using $param\_to\_kpis$.
9:     if common KPIs exist then
10:        Add an edge between $p_{i}$ and $p_{j}$ in $G$.
11:        Label the edge with the common KPIs.
12:  for each KPI $k$ do
13:     Find all ICPs associated with $k$.
14:     Group these ICPs in $P_{k}^{G}$.
15:  return  $P_{k}^{G}$.
When the conflict is detected, the CDC informs the CMC about the specific types of conflict occurring over the specific conflicting ICP. For the analysis of this paper, we adopt the priority conflict mitigation method proposed by Adamczyk et al. [1] and QoS-Aware Conflict Mitigation (QACM) method proposed in [3] to be deployed in the CMC. The priority method simply prioritises one conflicting xApp over the other based on their significance set by the MNO. On the other hand, the QACM considers specific QoS requirements of the conflicting KPIs and suggests an optimal value of the conflicting parameter at which the KPIs are either above their QoS mandates or as closer as possible to it. The following section discusses the simulation setup and results based on these two mitigation methods.
Figure 5: Network layout of the simulated O-RAN.
IV Experiment and Results
For the experimental analysis, we used MATLAB software for simulation, along with its 5G Toolbox and O-RAN 7.2 split [12, 3]. The simulation setup included two Next-generation NodeBs and ten User Equipments, as shown in Fig. 5. We focused on a direct conflict between the MRO xApp and ES xApp over Tx Power (TXP) values. The simulation parameters were: a Reference Signal Received Power (RSRP) threshold of -110 dBm for handover, a frequency of 2.4 GHz, and a simulation time of 10 minutes with a time step of 100 ms. Input control parameters included a default TXP value of 30 dBm, Cell Individual Offset (CIO) of 2 dB, Handover Hysteresis (HYS) of 0.5 dB, Time-To-Trigger (TTT) of 0.1 ms, Radio Electrical Tilt (RET) of 1.5 degrees, and an adjustment interval of 1000 ms. The UEs moved back and forth between the two gNBs with randomly assigned velocities between 0 and 5 m/s. The overall system goal is to minimize link failures and ping-pong handovers for MRO while maximizing energy efficiency for ES. A direct conflict arose when ES set the TXP to 3 dBm for energy savings, followed by MRO setting it to 50 dBm in the next request. We consider 5 conflict mitigation Methods, including a no-mitigation approach (NC) where the TXP was set to the recently requested value; a set back to default (SBD) method where the TXP was reset to its default value; a priority-based method where either ES (P-ES) or MRO (P-MRO) was prioritized; and finaly the QACM method was used to balance the TXP based on QoS thresholds. We repeated the simulation 500 times and the results were presented using box plots for each KPI.
(a) Energy efficiency.
(b) Number of link failures.
(c) Number of ping-pong handover.
Figure 6: Simulation results of KPIs adopting different conflict mitigation methods.
The Fig. 6(a) compares the energy efficiency across different conflict mitigation methods, including NC (No Conflict Mitigation), SBD (Set Back to Default), P-ES (prioritized ES), P-MRO (prioritized MRO), and QACM. The QACM method shows the highest energy efficiency (in bits/Joule), suggesting it is the most effective approach for maximizing energy savings, followed by P-ES. The NC, SBD and P-MRO methods exhibit lower energy efficiency indicating that prioritizing energy-saving strategies improves this KPI.
The Fig. 6(b) illustrates the impact of different conflict mitigation methods on reducing link failures. The SDB displays higher numbers of link failures followed by the NC and P-ES, implying poorer performance in maintaining stable connections. On the other hand, P-MRO and QACM show significantly fewer link failures with QACM being the most effective one. This indicates that prioritizing mobility and a balanced conflict resolution approach reduces the occurrence of link failures.
Finally, the Fig. 6(c) demonstrates the number of ping-pong handovers across the various methods. NC, SBD and P-ES have the highest frequency of ping-pong handovers, which may negatively affect network stability. In contrast, P-MRO and QACM exhibit significantly fewer ping-pong handovers with QACM providing the best mitigation. This shows that a mobility-optimized and balanced conflict resolution strategy can effectively reduce these undesirable handovers.
V Conclusion
In this paper, we provided an in-depth examination of xApp-level conflicts in O-RAN focusing on conflict detection and mitigation strategies. Through theoretical modeling and the use of conflict graphs, we identified different types of conflicts and highlighted their impacts on network KPIs. Our simulation results demonstrated the effectiveness of various mitigation methods, particularly the QACM approach, in managing conflicts while maintaining QoS thresholds. The results showed that QACM outperforms other mitigation strategies, including no conflict mitigation, reset to default and priority-based approaches in terms of energy efficiency and minimizing link failures & ping-pong handovers. Our future work will focus on refining conflict detection mechanisms experimenting with the detection mechanism in large-scale, exploring machine learning-based mitigation techniques, and evaluating real-world O-RAN deployments to further enhance the robustness of conflict management in dynamic network environments.
Acknowledgment
This research was partially funded by the European Union’s Horizon Europe research and innovation program through the Marie Skłodowska-Curie SE grant, under agreement number RE-ROUTE No 101086343.